OpenStudy (anonymous):

What is the correct simplification of the expression d^3 • d^5? In the form of multipulcation.

OpenStudy (anonymous):

When you have a common base (in this case, d) you simply add the exponents.

OpenStudy (anonymous):

but its multiplying...so i thought I multiplied the exponents?

OpenStudy (anonymous):

Well, you're multiplying two different things - the general rule for this is that if you're multiplying two exponential terms that have a common base, you add their exponents.

OpenStudy (anonymous):

In this case, you'd get d^3 * d^5 = d^8.

OpenStudy (anonymous):

Remember: \[a^m \times a^n = a^{m+n}\] And, \[(a^m)^n = a^{mn}\]

OpenStudy (anonymous):

hmmm..you sure hun?

OpenStudy (anonymous):

d^8.......note,,in multiplication with exponent,you must add the exponent of the two..

OpenStudy (anonymous):

hmm okay thx :) Im glad i know that now.

OpenStudy (anonymous):